4) A risk-averse consumer with $100,000 in wealth faces 0.1 probability of losing half of his wealth within the next year. a.NEED HELP!!!!!PLZ!
Answer #1

A) when lost half of wealth, then bad state wealth = 100,000/2

= 50,000

So Expected wealth = .9*100,000+ .1*50,000

= 90,000+5,000

= 95,000

b)

as when an insurance firm breaks even

then, net expected loss = net expected gain

πM = (1-π)*(B-M)

π: probability of good state

M : premium, B: promised benefit in bad state

πM = B-M - πB + πM

M = B-πB

M = B*(1-π)

B = 50,000

since full insurance, so entire loss is compensated in bad state

π= .9

so premium = .1*50,000

= 5000

C) now Maximum premium , willing to pay

= Initial wealth - Certainty Equivalent

= 100,000-85,000

= $ 15,000

​​​​​